QUESTION 1:

Nine randomly selected customers at a local fast-food restaurant ordered meals having the following calorie counts: 975, 520, 1560, 872, 1028, 533, 362, 954, and 1016. Let y denote the calorie content of a meal ordered at this restaurant. Find the following sum.

∑y2= Enter your answer in accordance to the question statement.

QUESTION 2:

The following data give the time (in minutes) that each of 20 students waited in line at their bookstore to pay for their textbooks in the beginning of Spring 2012 semester.

14 35 23 23 12
46 41 51 22 37
22 34 14 46 32
20 41 36 27 25

Construct a ranked stem-and-leaf display for the data.

QUESTION 3:

The 2011 gross sales of all firms in a large city have a mean of $2.4 million and a standard deviation of $ 0.6 million. Using Chebyshev's theorem, find at least what percentage of firms in this city had 2011 gross sales of $1.1 to $3.7 million.

Round your answer to the nearest whole number.

QUESTION 4:

A sample of 3000 observations has a mean of 86 and a standard deviation of 10. The distribution is bell-shaped. Using the empirical rule, find what percentage of the observations fall in the intervals x¯±1s, x¯±2s, and x¯±3s.


Enter the exact answers.


Approximately Enter you answer; 1s % of the observations fall in the interval x¯±1s.


Approximately Enter you answer; 2s % of the observations fall in the interval x¯±2s.


Approximately Enter you answer; 3s % of the observations fall in the interval x¯±3s.

QUESTION 5:

The following table gives the frequency distribution of the number of hours spent last week on cell phones (making phone calls and texting) by all 100 students of the tenth grade at a school.

Hours per Week

Number of Students

0 to less than 4

12

4 to less than 8

17

8 to less than 12

23

12 to less than 16

15

16 to less than 20

18

20 to less than 24

15


Find the mean, variance, and standard deviation.

Round your answers to two decimal places.

Mean = Enter you answer; Mean hours

Variance = Enter you answer; Variance

Standard deviation = Enter you answer; Standard deviation hours

QUESTION 7:

Consider the following two data sets.

Data Set I:                   12⁢    25⁢     37⁢    8⁢     41

Data Set II:                 24⁢    37⁢    49⁢    20⁢    53

Note that each value of the second data set is obtained by adding 12 to the corresponding value of the first data set. Calculate the standard deviation for each of these two data sets using the formula for sample data.



Round your answers to two decimal places.

Standard deviation of Data Set I = Enter you answer; Standard deviation of Data Set I

Standard deviation of Data Set II = Enter you answer; Standard deviation of Data Set II

QUESTION 8:

Consider the following two data sets.

Data Set I:                   12⁢    25⁢     37⁢    8⁢     41

Data Set II:                 24⁢    37⁢    49⁢    20⁢    53

Note that each value of the second data set is obtained by adding 12 to the corresponding value of the first data set. Calculate the standard deviation for each of these two data sets using the formula for sample data.



Round your answers to two decimal places.

Standard deviation of Data Set I = Enter you answer; Standard deviation of Data Set I

Standard deviation of Data Set II = Enter you answer; Standard deviation of Data Set II

QUESTION 9:

Following are the temperatures (in   ∘F) observed during eight wintry days in a midwestern city (sample):

                            23⁢    17⁢    6⁢    −7⁢    −2⁢    11⁢    16⁢    12

Find the range, variance, and standard deviation.



Round your answers to two decimal places, where appropriate.

Range = Enter you answer; Range   ∘F

Variance = Enter you answer; Variance

Standard deviation = Enter you answer; Standard deviation    ∘ F
Answer #1

"As per the Chegg guidelines, we need to answer for the first question only in multiple posting. Please ask single question in single post."

Solution:

Question 1:

Nine randomly selected customers at a local fast-food restaurant ordered meals having the following calorie counts:

Calorie content(y)
975
520
1560
872
1028
533
362
954
1016

Let us find y2 for the given data.

Calorie content(y) y2
975 975*975=950625
520 520*520=270400
1560 1560*1560=2433600
872 872*872=760384
1028 1028*1028=1056784
533 533*533=284089
362 362*362=131044
954 954*954=910116
1016 1016*1016=1032256
y2=7829298

First we need to square all the 'y' values and then add the results of 'y2' to get sum of 'y2' i.e..y2

Therefore the following sum is ∑y2=7829298